OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. X(0, 6), Y(5, 6) plzz help

OpenStudy (noorik.):

m = (y1-y2)/(x1-x2) m = (5-6)/(0-6) = -1/-6 then m = (y-y1)/(x-x1) 1/6 = (y-5)/(x-0) 6y-30 = x y = x/6 +5

OpenStudy (anonymous):

thanks i kinda under stand it

OpenStudy (anonymous):

what about this one Indicate in standard form the equation of the line passing through the given points.S(.5, 1), T(.5, 4)